OpenStudy (anonymous):

(5x-4) raised to the power of 1/2 = 6 answers: x=2 x=4 x=8 none

OpenStudy (lgbasallote):

\[(5x-4)^{1/2} = 6\] \[[(5x-4)^{\cancel{1/2}}]^{\cancel{2}} = 6^2\] \[5x - 4 = 6^2\] \[5x - 4 = 36\] can you solve for x now?

OpenStudy (anonymous):

x=8

OpenStudy (lgbasallote):

yup nice